# A mechanism to generate varying speed of light via Higgs-dilaton coupling:Theory and cosmological applications

### v2

@@ VI.2 Analyzing Pantheon Catalog of SNeIa using VSL
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{\mu_{i}^{\text{model}}-\mu_{i}^{\text{Pantheon}}}{\sigma_{i}^{\text{Pantheon}}}\biggr{)}^{2}\,.$ (62)
@@
VI.3 A reduced value of $H_{0}$ by a factor of $3/2$
@@ VI.5 Resolving the age problem, bypassing DE
$t_{0}^{\Lambda\text{CDM}}=\frac{2}{3\sqrt{\Omega_{\Lambda}}H_{0}}\,\text{arcsinh}\sqrt{\frac{\Omega_{\Lambda}}{\Omega_{M}}}\,.$ (64)
With $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, it yields an age of $13.6$ Gy, an accepted figure in standard cosmology. If dark energy is absent ($\Omega_{\Lambda}=0,\,\Omega_{M}=1$), Eq. (64) simplifies to $t_{0}^{\text{EdS}}=2/(3H_{0})$ which then, with $H_{0}\sim 71$, would result in an age of $9.2$ Gy which would be too short to accommodate the existence of the oldest stars—a paradox commonly referred to as the age problem.
@@ VI.6 Obtaining late-time acceleration via VSL, bypassing DE
It is an empirical fact that high-$z$ SNeIa’s appear fainter than what would be predicted by the EdS model [140, 141]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig. 1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
@@
VI.7 Physical intuition of late-time acceleration via VSL
@@
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic space as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SNeIa without the need for dark energy. This approach overcomes the fine-tuning and coincidence problems associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [146].
VI.8 A potential resolution to the $H_{0}$ tension via an astronomical origin
@@
$H_{0}(z)=H_{0}\,F(z)\,\biggl{(}1-\frac{3}{2}\frac{\ln F(z)}{\ln(1+z)}\biggr{)}^{-1}\,.$ (71)
@@
VI.9 An infinite cosmological horizon
@@
$l_{H}\left(t_{rec}\right)$ $=c_{0}\,t_{0}\,\left(\frac{t_{rec}}{t_{0}}\right)^{2/3}\int_{0}^{t_{rec}}\frac{dt^{\prime}}{t^{\prime}}=\infty\,.$ (73)
@@
VII Conclusions
The action given in Eqs. (7), (8), and (12) with $V(\Phi)\propto\Phi^{4}$ describes a generally covariant theory of gravitation and matter, which also possesses dilatation symmetry. This action is able to evade the observational bounds on the fifth force. A notable feature is the non-minimal coupling of matter with gravity, expressed through the coupling of the Higgs field $\Phi$ with a background dilaton field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. Crucially, the (dimensionful) Planck constant $\hbar$ and speed of light $c$ are absent in this adimensional action. Previously, Fujii and Wetterich utilized this action to generate variable particle masses and Newton’s ‘constant’ [12, 13, 14, 15, 16]. In Sections II and III of this paper, we developed an alternative mechanism to construct (variable) $\hbar$ and $c$ from the background dilaton field $\chi$. Our approach can be generalized when the matter Lagrangian (8) is extended to cover the Glashow–Weinberg–Salam (GWS) model of the $SU(3)\times SU(2)\times U(1)$ gauge group.
@@ VII Conclusions
In our approach, each open set vicinity of any given point $x^{*}$ on the spacetime manifold is equipped with a replica of the GWS action for matter fields (fermions, gauge vector bosons, and Higgs scalar bosons), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry holds locally within each open set. These GWS replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background dilaton field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. Crucially, within each open set, the length and time scales of physical processes are governed by the dilaton field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$, resulting in an anomalous anisotropic relation between the rate of clocks and the length of rods, $\tau^{-1}\propto l^{-3/2}$. This novel relation enables us to predict a new “time dilation effect of the Third kind”, to be reported elsewhere [154].
Our construction hence fulfills Einstein’s original—albeit long overlooked—vision in incorporating the local validity of Lorentz symmetry, the variation of speed of light in curved spacetimes, and the geometric nature of gravity in a unified framework [21].
The anisotropic relation $\tau^{-1}\propto l^{-3/2}$ mentioned above gives rise to new physics and phenomenology. Section VI of our paper presented key findings on the phenomenological consequences of our mechanism in the context of cosmology, particularly addressing late-time cosmic acceleration via the variation in $c$ while eliminating the need for dark energy. Remarkably, our results revive a crucial—but previously overlooked—discovery made by Blanchard, Douspis, Rowan-Robinson, and Sarkar in 2003 which circumvented the requirement for dark energy in their analysis of the CMB power spectrum [146].
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis and Eoin Ó Colgáin for their constructive and supportive comments during the development of this work.
@@
Appendix:  On Planck’s and Hartree’s unit systems
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The dilaton field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the dilaton field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary.

### v3

@@ VI.2 Analyzing Pantheon Catalog of SNeIa using VSL
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{\mu_{i}^{\text{model}}-\mu_{i}^{\text{Pantheon}}}{\sigma_{i}^{\text{Pantheon}}}\biggr{)}^{2}$ (62)
@@
VI.3 A reduced value of H0 by a factor of $3/2$
@@ VI.5 Resolving the age problem, bypassing DE
$t_{0}^{\Lambda\text{CDM}}=\frac{2}{3\sqrt{\Omega_{\Lambda}}H_{0}}\,\text{arcsinh}\sqrt{\frac{\Omega_{\Lambda}}{\Omega_{M}}}$ (64)
With $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, it yields an age of $13.6$ billion years, an accepted figure in standard cosmology. If dark energy is absent ($\Omega_{\Lambda}=0,\,\Omega_{M}=1$), Eq. (64) simplifies to $t_{0}^{\text{EdS}}=2/(3H_{0})$ which then, with $H_{0}\sim 71$, would result in an age of $9.2$ Gy which would be too short to accommodate the existence of the oldest stars—a paradox commonly referred to as the age problem.
@@ VI.6 Obtaining late-time acceleration via VSL, bypassing DE
It is an empirical fact that high-$z$ SNeIa’s appear fainter than what would be predicted by the EdS model [141, 142]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig.  1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
@@
VI.7 Physical intuition of late-time accelerationvia VSL, bypassing DE
@@
An alternative explanation for late-time acceleration is a decrease in the speed of light in the intergalactic space as the universe expands. In summary, VSL cosmology can account—qualitatively and quantitatively—for the distance modulus excess observed in high-$z$ SNeIa without the need for dark energy. This approach overcomes the coincidence problem associated with the concordance $\Lambda$CDM model. Furthermore, it bypasses the enigmatic nature of dark energy and aligns with BDRS’s 2003 groundbreaking analysis of the CMB power spectrum which also negates the dark energy hypothesis [147].
VI.8 H0 tension:   A potential resolution with an astronomical origin
@@
$H_{0}(z)=H_{0}\,F(z)\,\biggl{(}1-\frac{3}{2}\frac{\ln F(z)}{\ln(1+z)}\biggr{)}^{-1}$ (71)
@@
VI.9 An infinite cosmological horizon using VSL
@@
$l_{H}\left(t_{rec}\right)$ $=c_{0}\,t_{0}\,\left(\frac{t_{rec}}{t_{0}}\right)^{2/3}\int_{0}^{t_{rec}}\frac{dt^{\prime}}{t^{\prime}}=\infty$ (73)
@@
VII Discussion and Summary
On scale symmetry—Symmetries plays a fundamental role in constructing modern physical theories, with notable examples including general covariance in GR and gauge invariance in the Glashow–Weinberg–Salam (GWS) model of particle physics. In his seminal 1995 report [155], Bardeen proposed employing scale symmetry as a viable alternative to supersymmetry to cure the naturalness problem of the Higgs mass. It is well established that the Higgs particles, as quanta of scalar fields, suffer a quadratic divergence in their loop amplitudes, which destabilizes their ultraviolet-limit behavior and requires fine tuning. Bardeen suggested that classical scale invariance could protect the Higgs mass from this divergence. Theories that impose scale invariance in the matter and gravity sectors have been actively pursued in both particle physics and gravitational physics [332, 298, 299, 300, 301, 274, 275, 276, 277, 228, 190, 189, 259, 247, 254, 327, 357, 353, 349, 258, 330, 162, 340, 333, 347, 348, 356, 311, 233, 260, 170, 307, 215, 238, 239, 249, 313, 303, 261, 278, 248, 335, 358, 337, 183, 309, 184, 186, 187, 185, 262, 263, 241, 328, 329, 235, 237, 234, 236, 346, 192, 191, 193, 194, 304, 217, 218, 219, 221, 222, 220, 314, 321, 176, 160, 202, 204, 207, 206, 201, 200, 203, 205, 280, 334, 316, 319, 326, 264, 265, 281, 282, 284, 283, 240, 250, 294, 230, 231, 285, 214, 352, 188, 227, 255, 266, 232, 286, 226, 224, 225, 315, 287, 310, 308, 267, 268, 177, 178, 180, 179, 181, 182, 288, 289, 290, 164, 169, 168, 167, 165, 166, 242, 243, 244, 245, 246, 223, 351, 251, 252, 253, 355, 197, 229, 302, 256, 171, 322, 323, 325, 324, 318, 345, 291, 216, 295, 163, 359, 211, 212, 213, 341, 279, 342, 343, 269, 336, 320, 270, 271, 312, 195, 196, 257, 317, 161, 172, 344, 199, 198, 272, 157, 156, 158, 159, 338, 339, 297, 354, 305, 306, 209, 210, 350, 292, 273, 331, 208, 173, 174, 175, 293, 296, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129]. Scale invariance typically necessitates (non-minimal) coupling of the Higgs field with gravity via a scalar field $\chi$—commonly referred to as a “dilaton”—which arises naturally in various theoretical contexts, such as Kaluza–Klein, string theory, and braneworld scenarios [12].
In Section II, we consider a scale-invariant action of gravity and matter, as described by Eqs. (7), (8), and (12) with $V(\chi)\propto\chi^{4}$. This action is generally covariant, locally Lorentz invariant, and $U(1)$ gauge invariant (for the matter sector). Notably, the scale invariance of the action allows it to evade the observational bounds on the fifth force [123, 115]. The Higgs field couples with the dilaton field $\chi$ in the form $\chi^{2}\,\Phi^{2}$. In addition, the matter Lagrangian $\mathcal{L}_{\text{mat}}$, given in Eq. (8) as a prototype, is invariant with respect to a discrete $\mathbb{Z}_{2}$ symmetry of the Higgs field, viz. $\Phi\leftrightarrow-\Phi$, although it can be readily generalized to encompass the $SU(3)\times SU(2)\times U(1)$ gauge group of the GWS model.
Our mechanism to generate variable $\hbar$ and $c$—Because of scale invariance, all parameters in the action are dimensionless. This crucial fact means that the Planck constant $\hbar$ and the speed of light $c$—as dimensionful quantities—must be absent from the outset of this adimensional action.
In Section III, we developed—for the first time—a recipe to construct $\hbar$ and $c$ from the Higgs-dilaton coupling in the matter sector $\mathcal{L}_{\text{mat}}$ of Eq. (8). The metric $g_{\mu\nu}$ and the dilaton $\chi$ are treated as slowly varying background fields. Within a local set surrounding a given point $x^{*}$, the discrete $\mathbb{Z}_{2}$ symmetry of the Higgs field is spontaneously broken due to the quartic “potential” $\frac{1}{2}\mu^{2}\,\chi^{2}\,\Phi^{2}-\frac{1}{4}\lambda\,\Phi^{4}$ (for $\lambda>0$). The Higgs field then acquires a non-zero VEV $\left\langle\Phi(x^{*})\right\rangle$ which is proportional to $\chi(x^{*})$; see Eq. (13). The non-zero Higgs VEV gives rise to mass for both the fermion and the Higgs scalar boson (while the gauge vector boson remains massless since the $U(1)$ gauge remains unbroken). Importantly, the Higgs VEV allows us to construct a Planck constant $\hbar_{*}$ and a speed of light $c_{*}$ to be effective within the local set around $x^{*}$.
It is worth noting that in Ref. [135], we presented a “short-cut” approach. Instead of invoking the Higgs-dilaton coupling $\chi^{2}\,\Phi^{2}$ and the Yukawa coupling between the fermion and the Higgs field $\bar{\psi}\psi\,\Phi$ as we do in Eq. (8) in this current paper, therein we intentionally suppressed the Higgs field and let the fermion directly couple with the dilaton via the term $\chi\,\bar{\psi}\psi$. In doing so, we also obtained a construction for $\hbar_{*}$ and $c_{*}$ similar to that presented in this paper. However, the current paper aims to explicitly incorporate the Higgs field and utilize the mechanism of spontaneous symmetry breaking. so that a connection to the GWS model for the matter sector can be established.
@@ VII Discussion and Summary
In either approach, the end result is that each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Quantum Electrodynamics (QED) action for matter fields (i.e. fermion, $U(1)$ gauge vector boson, and Higgs scalar boson), operating with its own effective values of $\hbar_{*}$ and $c_{*}$. Lorentz symmetry is valid locally within each open set. These QED replicas are isomorphic (i.e. structurally identical), differing only in the values of $\hbar_{*}$ and $c_{*}$, which are determined by the prevailing value of the background dilaton field $\chi(x^{*})$ per $\hbar_{*}=\chi^{-1/2}(x^{*})$ and $c_{*}=\chi^{1/2}(x^{*})$. With the gravitational Lagrangian $\mathcal{L}_{\text{grav}}$ in Eq. (12) governing the dynamics of $\chi$, the values of $\hbar_{*}$ and $c_{*}$ thus vary from one open set to another across spacetime.
Crucially, within each open set, the length and time scales of physical processes are governed by the dilaton field as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relationship between the rate of clocks and the length of measuring rods, per $\tau^{-1}\propto l^{-3/2}$. This novel relationship allows us to predict a new “time dilation effect of the Third kind”, to be reported in future work.
The role of VSL in late-time cosmologyThe anisotropic relationship $\tau^{-1}\propto l^{-3/2}$ has immediate cosmological consequences. Section VI presents key findings regarding the impacts of variable speed of light (VSL) on the Hubble diagram of SNeIa, with full technical details of our analysis provided in a companion paper, Ref. [21].
In the presence of VSL, the standard cosmography is no longer applicable. Section VI.1 produces a new cosmography that accommodates VSL:
(i) Under the assumption that the dilaton is related to the cosmic scale factor as $\chi\propto a^{-1}$, the FLRW metric is modified to
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\left[dr^{2}+r^{2}d\Omega^{2}\right]$
which leads to a modified Lemaître redshift formula (where the function $F(z)$ accounts for variations in the dilaton value among individual galaxies)
$1+z=a^{-3/2}\,F^{3/2}(z)$
and a modified Hubble law
$z=\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}$
[…]
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}$
Based on this formula, in Section VI.2, we (re)analyzed the Pantheon Catalog of SNeIa. The key findings are as follows:
* We achieved a fit to the Pantheon data that surpasses the quality of the $\Lambda$CDM model. The function $F(z)$ is parametrized as $F(z)=1-(1-F_{\infty}).\left(1-(1+z)^{-2}\right)^{2}$ with $F_{\infty}=0.931\pm 0.11$ (95% CL).
* We obtained a new Hubble value $H_{0}=47.2\pm 0.4$ (95% CL), which is approximately $2/3$ of the $H_{0}\approx 70$ value derived from SNeIa when relying on the $\Lambda$CDM model (with $\Omega_{\Lambda}\approx 0.7$). The reduction in the $H_{0}$ value arises from the $3/2$–multiplicative factor in the modified Hubble law, Eq. (57). See Section VI.3.
* The reduced value of $H_{0}=47.2$ yields a value of $t_{0}=13.9\pm 0.1$ Gy (95% CL) for the age of an EdS universe, effectively resolving the age paradox without requiring dark energy. See Section VI.5.
* Additionally, the function $F(z)$ offers a potential resolution of the $H_{0}$ tension. See Section VI.8.
Importantly, Sections VI.6 and VI.7 provide a new perspective: the Hubble diagram of SNeIa does not necessarily indicate an accelerating universe. Instead, it supports an alternative interpretation in favor of an EdS universe characterized by a declining speed of light.
On the Blanchard–Douspis–Rowan-Robinson–Sarkar (BDRS) analysis of the CMB power spectrum, avoiding dark energy—Section VI.4 revisits an important—but long-overlooked—proposal from 2003 [147], where BDRS shed new light on the CMB power spectrum. By adopting a double-power form for the primordial fluctuation spectrum as given Eq.  (63), they achieved an excellent fit to WMAP’s CMB power spectrum within an EdS model, namely, without invoking dark energy. A surprising outcome of their analysis was a value of $H_{0}\approx 46$, which, while at odds with the canonical range of $H_{0}\sim 67-73$, aligns remarkably well with the $H_{0}=47.2$ value derived from our analysis of SNeIa. The nearly perfect agreement in the values of $H_{0}$, obtained from two datasets with different natures and covering separate epochs, indicates a consistent cosmological framework based on the EdS model with a variable speed of light, while eliminating the need for dark energy.
VIII Conclusion
From the Higgs-dilaton coupling, we have derived variations in Planck’s quantum of action and the speed of light in curved spacetime. Our work hence—for the first time—fulfills Einstein’s original vision, dating back to 1911-1912 [23, 24, 25], by integrating the local validity of Lorentz symmetry, the variability of the speed of light, and the (Riemannian) geometric nature of gravity into a unified framework.
Importantly, the variable speed of light provides an alternative explanation for the Hubble diagram of SNeIa: Rather than being attributed to a late-time cosmic acceleration, the excess in the distance modulus observed in high-redshift SNeIa is a result of a declining speed of light in an expanding Einstein–de Sitter universe.
Strikingly, the reduced value of $H_{0}=47.2$ derived herein from the SNeIa data aligns remarkably well with the $H_{0}\approx 46$ value that BDRS deduced from the CMB data, all while bypassing the need for dark energy [147]. Taken together, our work and that of BDRS diminish the necessity for the Dark Energy hypothesis, along with its associated fine-tuning and coincidence problems.
In addition, potential implications of scale invariance and our mechanism in realms beyond late-time cosmology are briefly outlined in Appendix A.
Lastly, as $\hbar$ and $c$ become variable in scale-invariant gravity, they can no longer serve as “fundamental units” within Planck’s $cGh$ system [7, 8, 9]. Appendix B introduces a new unit system suitable for variable $\hbar$ and $c$.
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis, Eoin Ó Colgáin, Subir Sarkar, Leandros Perivolaropoulos, Demosthenes Kazanas, John Moffat, Alberto Salvio, Ilya L. Shapiro, Gregory Volovik and Bruce Bassett for their constructive and insightful feedback. I wish to thank the anonymous Referee for his or her comments, which significantly helped improve the paper.
@@
Appendix A   AN OUTLOOK BEYOND LATE-TIME COSMOLOGY
Here, we speculate on the potential ramifications of scale invariance, as well as variable $c$ and $\hbar$, in realms beyond late-time cosmology.
* Early-time cosmology: Section VI.9 “An infinite cosmological horizon using VSL” explicitly shows that an EdS universe with declining speed of light exhibits an infinite cosmological horizon; see Eq. (73). This result could help resolve the horizon paradox, which was one of the key motivations for the VSL proposals made by Moffat and the Albrecht–Magueijo team in the 1990s [26, 27].
* On the flattening galactic rotation curves: For $\mathcal{L}_{\text{grav}}$ in Eq. (12), the case of $\omega=0$ and $V(\chi)\propto\chi^{4}$ corresponds to quadratic gravity. It has been shown [128, 124] that when expanding around a de Sitter background metric (rather than the flat background conventionally used), the massless rank-2 tensor graviton mode produces a long-range gravitational potential with the correct Newtonian tail, viz. $\sim 1/r$. Furthermore, it has been shown [124] that the massless rank-0 scalar graviton mode yields a long-range potential which grows linearly with distance, viz. $\sim r$. A superposition of these modes could give rise to the Mannheim–Kazanas potential (previously derived for conformal gravity [360, 361]), $-GM/r+\gamma\,r$, where $GM$ and $\gamma$ are two adjustable parameters. In [363, 362], this potential has been employed to successfully account for the observed flattening of rotation curves across a wide range of galaxies, without invoking dark matter.
* Variable $\hbar$: A Planck constant that varies in spacetime could, in principle, alter—both qualitatively and quantitatively—the radiative behavior of black holes in scale-invariant gravity.
* On the Hierarchy Problem and the Cosmological Constant Problem: In quantum field theories of matter in flat spacetime, the loop corrections to the Higgs mass and vacuum energy manifest as quadratic and quartic divergences, respectively [366, 367, 365, 364, 368]. However, it is plausible to suspect that these divergences only appear problematic when considered in the context of flat spacetime—namely, in isolation from the otherwise curved background spacetime. That is to say, these loop corrections currently account for contributions solely from the matter sector, while neglecting those from the gravitational sector. One intuitive way to understand this is that in scale-invariant gravity, as $\hbar$ and $c$ become dependent on the dilaton per $\hbar\propto\chi^{-1/2}$ and $c\propto\chi^{1/2}$, the dilaton—as a component of the gravitational sector—could, in turn, modulate these loop corrections and keep them in check. If so, scale invariance, as inspired by Bardeen’s seminal idea [155], might open a pathway toward protecting the Higgs mass and vacuum energy from divergence and fine tuning.
These intriguing possibilities, albeit speculative at this stage, represent open avenues for future investigation.
Appendix B   ON PLANCK’S AND HARTREE’S UNIT SYSTEMS
In our mechanism, the electron mass $m$ and charge $e$ are dimensionless quantities derived from the parameters $f$, $\mu$, $\lambda$, and $\alpha$ via Eq. (19). The dilaton field $\chi$ is the single entity that carries a scale, which then serves as the only unit available. Consequently, all physical quantities can be expressed exclusively in terms of the magnitude of $\chi$. For instance, $\hbar$ and $c$ have the units of $\chi^{-1/2}$ and $\chi^{1/2}$, respectively, whereas $G$ is dimensionless. The rest mass energy $E=mc^{2}$ has the unit of $\chi$, owing to $m$ being dimensionless and $c\propto\chi^{1/2}$. Similarly, the Coulomb potential of a point charge $e$, given by $e^{2}/r$, also carries the unit of $\chi$, by virtue of $e$ being dimensionless and $r\propto\chi^{-1}$. In general, a theory with dilatation symmetry would require only one ‘unit’ (associated with the dilaton field $\chi$), rendering the canonical three-unit system (mass $M$, length $L$, time $T$) unnecessary. It is worth noting that this conclusion agrees with Ref. [369], which deduced from a spacetime perspective that the number of fundamental constants equals one in relativistic spacetimes.